Suppose the population standard deviation is known. Suppose we take three independent samples and report a 95% confidence interval for the population mean μ for each set of data. What is the probability that a majority of the three confidence intervals contain the true population mean μ (rounded to four decimal places)? (Remember to express your answer as a proportion.)
Answer #1
here this is binomial with parameter n=3 and p=0.95

probability that a majority of the three confidence intervals contain the true population mean μ

=P(X=2)+P(X=3)=(3C2)*(0.95)2*(0.05)1+(3C3)*(0.95)3*(0.05)0 =3971/4000 =0.9928